OpenStudy (anonymous):

Box A has a volume of 48 cubic meters. Box B is similar to box A. To create box B, box A's dimensions were doubled. What is the volume of box B? (1 point) Select one: a. 24 m3 b. 96 m3 c. 288 m3 d. 384 m3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@prowrestler do you have any ideas how to start this problem off?

OpenStudy (anonymous):

no thats why i posted it

OpenStudy (anonymous):

Okay, well we know to calculate volume of box A that the formula is: \(V_a = l\times w\times h\) Do you see where I am going?

OpenStudy (anonymous):

Yes i do

OpenStudy (anonymous):

Recall that the question says: "To create box B, box A's dimensions were doubled." \[V_b=l(2)\times w(2) \times h(2)\] You could also shorten this by: \[V_b=8(l\times w \times h)\] Do you know what to do next?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

What are you confused about?

OpenStudy (anonymous):

u multiply the volume by eight is what languageenthusiast is saying

OpenStudy (anonymous):

@TazMann Is correct.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thnx for the help

OpenStudy (anonymous):

np anytime